# The effect of priors on Learning with Restricted Boltzmann Machines

### v1

@@ 4 Exploring arbitrary priors: Mismatched Setting
Transition to the Spin Glass Phase
@@
$P_{\Omega_{\xi}}(\xi)\left(1+\sqrt{\alpha\beta q_{\tau}}z\xi+O(q_{\tau})\right)e^{\frac{\alpha}{2}\beta(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}}$ (40)
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta q}z\tau+O(q)\right)e^{\frac{\beta}{2}d\,\tau^{2}}\,.$ (41)
@@
$d_{\tau}$ $=$ $\langle\tau^{2}\rangle_{0}=\mathbb{E}_{\tau}[\tau^{2}e^{\frac{\beta}{2}d\,\tau^{2}}]/\mathbb{E}_{\tau}[e^{\frac{\beta}{2}d\,\tau^{2}}],\ \ \ \ \ \ \tau\sim P_{\Omega_{\tau}}$ (42)
$d$ $=$ $\langle\xi^{2}\rangle_{0}=\mathbb{E}_{\xi}[\xi^{2}e^{\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}}]/\mathbb{E}_{\xi}[e^{\frac{\alpha}{2}\beta(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}}],\ \ \ \xi\sim P_{\Omega_{\xi}}\,,$ (43)
@@
From Eqs. (44-45) a second order phase transition to the spin glass phase may occur at
@@
that can be solved to obtain the critical temperature $T_{c}(\alpha)$ or equivalently the critical size $\alpha_{c}(T)$. Note that there is no dependence of $\hat{\beta}$, meaning that the spin glass transition only depends on a (too low) inference temperature of the student. As expected [18, 7, 51] the P-SG transition (46) corresponds to the P-F transition in the Nishimori regime (39). Moreover it also corresponds to the P-SG transition of a generalized Hopfield model with random patterns at inverse temperature $\beta$ and load $\alpha$. Indeed, the expression (46) matches the one found in [19]. For example fixing $\Omega_{\xi}=\Omega_{\tau}=0$ one gets the bipartite SK P-SG transition $T_{c}(\alpha)=\sqrt{\alpha}$, while using $\Omega_{\xi}=0$ and $\Omega_{\tau}=1$ it is possible to recover the P-SG transition line of the Hopfield model $T_{c}(\alpha)=1+\sqrt{\alpha}$.
Transition to the Retrieval Phase
When the size of the dataset is sufficiently large, lowering the temperature may lead to a continuous transition towards a signal retrieval (sR) region. The transition line can be obtained by expanding Eq. (14) for small $m$ and $m_{\tau}$, keeping $p=p_{\tau}=q=q_{\tau}=0$. As previously done, first of all Eqs. (17-18) can be expanded as
$P_{\Omega_{\xi}}(\xi)\left(1+\alpha\sqrt{\beta\hat{\beta}}m_{\tau}\hat{\xi}\xi+O(m_{\tau})\right)e^{\frac{\alpha}{2}\beta(\,d_{\tau}-\langle\tau^{2}\rangle_{\tau}\,)\xi^{2}}$ (47)
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta\hat{\beta}}m\tau\hat{\tau}+O(m)\right)e^{\frac{\beta}{2}d\tau^{2}},$ (48)
@@
$m$ $=$ $\big{\langle}\hat{\xi}\langle\xi\rangle_{\xi|z,s,\hat{\xi}}\big{\rangle}_{z,s,\hat{\xi}}=\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\langle\hat{\xi}^{2}\rangle_{\hat{\xi}}\langle\xi^{2}\rangle_{0}+o(m_{\tau})=\alpha\sqrt{\hat{\beta}\beta}\,m_{\tau}\,d\,\langle\hat{\xi}\rangle_{\hat{\xi}}+o(m_{\tau})$ (49)
$m_{\tau}$ $=$ $\big{\langle}\hat{\tau}\langle\tau\rangle_{\tau|z,s,\hat{\tau}}\big{\rangle}_{z,s,\hat{\tau}}=\sqrt{\hat{\beta}\beta}m\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}\langle\tau^{2}\rangle_{0}+o(m)=\sqrt{\hat{\beta}\beta}\,m\,d_{\tau}\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}+o(m),$ (50)
where again $\langle\cdot\rangle_{0}$ denotes the expectations w.r.t. the distributions of Eqs. (17) or (18) at zero effective field. In Eq.(49) we can use $d_{\hat{\xi}}:=\langle\hat{\xi}^{2}\rangle_{\hat{\xi}}=1$ which derives from the analysis of the teacher partition function (89). Again the self-overlaps $d$ and $d_{\tau}$ are solutions of Eq.(42-43), which implies $d=1$. Recalling that .$\left\langle.\right\rangle_{\hat{\tau}}$ is the average over the tilted distribution $P_{\Omega_{\hat{\tau}}}(\hat{\tau})e^{\frac{\hat{\beta}}{2}\hat{\tau}^{2}}$, it holds
$d_{\hat{\tau}}$ ::=:= $\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}=\frac{1-\hat{\beta}\Omega_{\hat{\tau}}^{2}}{(1-\hat{\beta}\Omega_{\hat{\tau}})^{2}}\,,$ (51)
thus, from Eqs. (49-50), a second order phase transition to the sR phase may occur at
@@
For example when both the T-RBM and S-RBM have only binary units the critical temperature reads as $T_{c}(\alpha)=\alpha\hat{\beta}$ while in the case of the classical Hopfield model, i.e. $\Omega_{\tau}=\Omega_{\hat{\tau}}=1$ [7], $T_{c}(\alpha)=1+\alpha\hat{\beta}\;.$ Since $d=1$ and $d_{\tau}$ doesn’t depend of $\alpha$, the P-sR critical temperature $T_{c}(\alpha)$ grows linearly with the dataset’s size $\alpha$. Conversely, from Eq. (46), the critical temperature of the P-SG transition grows only as $O(\sqrt{\alpha})$. This means that for larger values of $\alpha$ the P-sR phase transition occurs first. On the contrary for small $\alpha$ the spin glass phase emerges at a higher temperature.
@@
$\begin{cases}1=\alpha\beta^{2}d_{\tau}^{2}\\ 1=\alpha\beta\hat{\beta}d_{\tau}d_{\hat{\tau}},\end{cases}$ (53)
@@
$P_{triple}=(\alpha_{c},T_{c})=\left(\frac{1}{\hat{\beta}^{2}d_{\hat{\tau}}^{2}}\ ,\ \frac{d_{\tau}}{\hat{\beta}d_{\hat{\tau}}}\right).$ (54)
@@
These properties can be observed from the phase diagrams obtained by different choices of the priors, which are presented in Figs. (4-8). All the depicted transition lines from the SG region to the sR or the eR regions can be found numerically by solving the saddle point Eqs. (13-16).
@@
Finally, the example retrieval region may emerge at low inference temperature when $\alpha$ is very close to zero. This phase corresponds to the pattern retrieval region in the generalized Hopfield models [19] under the critical capacity. In this region the student pattern tends directly to be aligned with one of the examples instead of learning their archetype, i.e. the teacher pattern. Unlike the sR regime, this is a phase where the machine is trying to learn by memorization and not by generalization [7].
The role of the S-RBM priors
@@
In Fig. (4) the phase diagrams of different S-RBMs with binary patterns ($\Omega_{\xi}=0$) are shown in the case of a binary T-RBM, i.e. $\Omega_{\hat{\xi}}=\Omega_{\hat{\tau}}=0$. One can observe the appearance of the eR phase at low temperature as soon as the student hidden unit starts to have a Gaussian tail, i.e. $\Omega_{\tau}>0$. The critical temperature at $\alpha=0$ is exactly $\Omega_{\tau}$. In the limit $\Omega_{\tau}\to 1$ this phase extends up to the critical capacity of the Hopfield model $\alpha=0.14$. As expected the critical size $\alpha_{c}$ of the triple point doesn’t change with the student priors while its temperature increases with $\Omega_{\tau}$. As a consequence the area of the sR phase expands in the direction of higher temperatures. This is typically an advantage for MC algorithms that samples from the posterior starting from high temperature like simulated annealing [11, 32].
@@
In Fig. (5), we observe similar behavior for a dataset generated by a Hopfield like T-RBM, i.e. $\Omega_{\hat{\tau}}=1$, $\Omega_{\hat{\xi}}=0$. The different generation procedures significantly affect the number of data points required for the student to transition into the sR regime: as prescribed by Eq.(54), $\alpha_{c}$ is smaller than that of Fig.(4). One can in principle investigate the effect of changing $\Omega_{\xi}$, however we have verified that the phase diagram is not affected by that prior since it only enters in the self-overlap, and $d=1$ is always solution of Eqs. (13-16). Interestingly, the benefit of using Gaussian units—where at any given temperature, the sR region occurs at smaller $\alpha$—remains consistent regardless of the dataset properties. Specifically, it is not always advantageous for the student to align their hyperparameters with those of the teacher.
The role of the T-RBM priors
@@
Figure 6: Effect of the teacher architecture ($\hat{\beta}=0.8$) on the student’s phase diagram. The signal’s prior is defined by $\Omega_{\hat{\xi}}=0$. The analyzed student is working with a Hopfield architecture: $\Omega_{\xi}=0$,$\Omega_{\tau}=1$. The more $\hat{\tau}$ is gaussian the more easily the student retrieves the pattern $\bm{\hat{\xi}}$, in particular $P_{\text{triple}}$ decreases in both its coordinates.
Fig. (6) shows the effect of changing $\Omega_{\hat{\tau}}$. It is possible to appreciate the fact that the information present in the data increases notably the more the $\hat{\tau}$ prior becomes Gaussian. This can be seen by the reduced amount of generated samples we need to enter the sR regime (a similar effect of increasing $\hat{\beta}$ in [7]). On the other hand, the critical temperature of the student at the triple point, $T_{c}$, is lowered. Nevertheless, if we fix an inference temperature, the student enters the sR regime more easily as $\Omega_{\hat{\tau}}\to 1$.
@@
Finally in Fig. (8) we explore the role of the dataset unit prior $\Omega_{s}$. The different values of $\Omega_{s}$ do not change the triple point $P_{c}$. As one can observe the only effect is the shift of the eR-SG transition line. This is related to the memorization capability of the S-RBM and is in agreement with the phase diagrams in [19].
Spherical regularization
In the previous sections we have seen that the term $z(\bm{\xi})^{-M}$ in the posterior plays the role of a regularization enforcing the Nishimori identity $d=1$. Without that regularization the model would be ill defined in the case of an architecture with both $\Omega_{\xi}$ and $\Omega_{\tau}$ different from zero. This is a typical problem with fully Gaussian disordered systems at low temperature [19, 16, 17, 31].
From a practical point of view is more convenient to replace that term directly with a spherical constraint $\delta\left(N-\sum_{i=1}^{N}\xi_{i}^{2}\right)$. The computation leading to the set of saddle point Eqs. (13-16) can be exactly followed with the only addition of a lagrange multiplier $\omega$ in the gaussian weight for the $\xi$ variable, i.e. its effective distribution reads as
$P_{\Omega_{\xi}}(\xi)e^{-\frac{\alpha\beta}{2}\left(\alpha\beta q_{\tau}-d_{\tau}+\omega\right)\xi^{2}+\xi\left(\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\hat{\xi}+\sqrt{\alpha\beta q_{\tau}}z+\beta\Omega_{\tau}ps\right)}\,,$ (55)
where $\omega$ has to be fixed to obtain $d=1$. Making a comparison with Eq.(17) one can see that the correct choice is $\omega=\langle\tau^{2}\rangle_{\tau}$. From the computational point of view it is more convenient to adjust the Lagrange multiplier during the solution of the saddle point equations. In Fig.(9) it is shown the phase diagram where the priors of both the weights and the student hidden variables have Gaussian tails. The weights have been regularized with a spherical constraint. One can compare these results with the right panel of Fig.(4), where $\Omega_{\xi}=0$. It is possible to see the presence of the eR phase, due to $\Omega_{\tau}=1$, that tends to quickly disappear as $\Omega_{\xi}\to 1$, in agreement with [19]. It is also observed a bending of the sR region at low temperature that however is not particularly significant since that is the region where RSB corrections are expected. Finally, as expected, the triple point position doesn’t change.
Figure 9: Spherical regularization effect on the retrieval capability of the model (case with $\Omega_{\tau}=1$). Using binary data from a binary T-RBM ($\hat{\beta}=0.8$), it is possible to see that $\alpha_{c}$ and $T_{c}$ do not move. The only effect from the changes of $\Omega_{\xi}$ is the compression of the $eR$ phase.
Conclusions

### v2

@@ 5 Exploring arbitrary priors: Mismatched Setting
5.1 Transition to the Spin Glass Phase
@@
$P_{\Omega_{\xi}}(\xi)\left(1+\sqrt{\alpha\beta q_{\tau}}z\xi+O(q_{\tau})\right)\exp\left(\frac{\alpha}{2}\beta(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}\right)$ (37)
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta q}z\tau+O(q)\right)\exp\left(\frac{\beta}{2}d\,\tau^{2}\right)\,.$ (38)
@@
$d_{\tau}=\langle\tau^{2}\rangle_{0}=\mathbb{E}_{\tau}[\tau^{2}\exp\left(\frac{\beta}{2}d\,\tau^{2}\right)]/\mathbb{E}_{\tau}[\exp\left(\frac{\beta}{2}d\,\tau^{2}\right)],\ \ \ \ \ \ \tau\sim P_{\Omega_{\tau}}$ (39)
$\hskip-42.67912ptd=\langle\xi^{2}\rangle_{0}=\mathbb{E}_{\xi}[\xi^{2}\exp\left(\frac{\alpha\beta}{2}(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}\right)]/\mathbb{E}_{\xi}[\exp\left(\frac{\alpha}{2}\beta(d_{\tau}-\langle\tau^{2}\rangle_{\tau})\xi^{2}\right)],\ \ \ \xi\sim P_{\Omega_{\xi}}\,$, (40)
@@
From Eqs. (41-42) a second order phase transition to the spin glass phase may occur at
@@
that can be solved to obtain the critical temperature $T_{c}(\alpha)$ or equivalently the critical size $\alpha_{c}(T)$. Note that there is no dependence of $\hat{\beta}$, meaning that the spin glass transition only depends on a (too low) inference temperature of the student. As expected barra2017phase , alemanno2023hopfield , theriault2024dense
the P-SG transition (43) corresponds to the P-F transition in the Nishimori regime (36). Moreover it also corresponds to the P-SG transition of a generalized Hopfield model with random patterns at inverse temperature $\beta$ and load $\alpha$. Indeed, the expression (43) matches the one found in barra2018phase .
For example fixing $\Omega_{\xi}=\Omega_{\tau}=0$ one gets the bipartite SK P-SG transition $T_{c}(\alpha)=\sqrt{\alpha}$, while using $\Omega_{\xi}=0$ and $\Omega_{\tau}=1$ it is possible to recover the P-SG transition line of the Hopfield model $T_{c}(\alpha)=1+\sqrt{\alpha}$.
5.2 Transition to the Retrieval Phase
When the size of the dataset is sufficiently large, lowering the temperature may lead to a continuous transition towards a signal retrieval (sR) region. The transition line can be obtained by expanding Eq. (11) for small $m$ and $m_{\tau}$, keeping $p=p_{\tau}=q=q_{\tau}=0$. As previously done, first of all Eqs. (14-15) can be expanded as
$P_{\Omega_{\xi}}(\xi)\left(1+\alpha\sqrt{\beta\hat{\beta}}m_{\tau}\hat{\xi}\xi+O(m_{\tau})\right)\exp\left(\frac{\alpha}{2}\beta(\,d_{\tau}-\langle\tau^{2}\rangle_{\tau}\,)\xi^{2}\right)$ (44)
$P_{\Omega_{\tau}}(\tau)\left(1+\sqrt{\beta\hat{\beta}}m\tau\hat{\tau}+O(m)\right)\exp\left(\frac{\beta}{2}d\tau^{2}\right),$ (45)
@@ 5.2 Transition to the Retrieval Phase
$\hskip-56.9055ptm$ $=\big\langle\hat{\xi}\langle\xi\rangle_{\xi|z,s,\hat{\xi}}\big\rangle_{z,s,\hat{\xi}}=\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\langle\hat{\xi}^{2}\rangle_{\hat{\xi}}\langle\xi^{2}\rangle_{0}+\mathcal{O}(m_{\tau})=\alpha\sqrt{\hat{\beta}\beta}\,m_{\tau}\,d\,\langle\hat{\xi}\rangle_{\hat{\xi}}+\mathcal{O}(m_{\tau})$ (46)
$m_{\tau}$ $=\big\langle\hat{\tau}\langle\tau\rangle_{\tau|z,s,\hat{\tau}}\big\rangle_{z,s,\hat{\tau}}=\sqrt{\hat{\beta}\beta}m\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}\langle\tau^{2}\rangle_{0}+\mathcal{O}(m)=\sqrt{\hat{\beta}\beta}\,m\,d_{\tau}\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}+\mathcal{O}(m),$ (47)
where again $\langle\cdot\rangle_{0}$ denotes the expectations w.r.t. the distributions of Eqs. (14) or (15) at zero effective field. In Eq.(46) we can use $d_{\hat{\xi}}:=\langle\hat{\xi}^{2}\rangle_{\hat{\xi}}=1$ which derives from the analysis of the teacher partition function (90). Again the self-overlaps $d$ and $d_{\tau}$ are solutions of Eq.(39-40), which implies $d=1$. Recalling that $\left\langle.\right\rangle_{\hat{\tau}}$ is the average over the tilted distribution $P_{\Omega_{\hat{\tau}}}(\hat{\tau})\exp\left(\frac{\hat{\beta}}{2}\hat{\tau}^{2}\right)$, it holds
$d_{\hat{\tau}}$ $:=$ $\langle\hat{\tau}^{2}\rangle_{\hat{\tau}}=\frac{1-\hat{\beta}\Omega_{\hat{\tau}}^{2}}{(1-\hat{\beta}\Omega_{\hat{\tau}})^{2}}\,$, (48)
thus, from Eqs. (46-47), a second order phase transition to the sR phase may occur at
@@
For example when both the T-RBM and S-RBM have only binary units the critical temperature reads as $T_{c}(\alpha)=\alpha\hat{\beta}$ while in the case of the classical Hopfield model, i.e. $\Omega_{\tau}=\Omega_{\hat{\tau}}=1$ alemanno2023hopfield , $T_{c}(\alpha)=1+\alpha\hat{\beta}\;$. Since $d=1$ and $d_{\tau}$ doesn’t depend of $\alpha$, the P-sR critical temperature $T_{c}(\alpha)$ grows linearly with the dataset’s size $\alpha$. Conversely, from Eq. (43), the critical temperature of the P-SG transition grows only as $O(\sqrt{\alpha})$. This means that for larger values of $\alpha$ the P-sR phase transition occurs first. On the contrary for small $\alpha$ the spin glass phase emerges at a higher temperature.
@@
[…]
@@
[…]
@@
These properties can be observed from the phase diagrams obtained by different choices of the priors, which are presented in Figs. (4-9). All the depicted transition lines from the SG region to the sR or the eR regions can be found numerically by solving the saddle point Eqs. (10-13).
@@
Finally, the example retrieval region may emerge at low inference temperature when $\alpha$ is very close to zero. This phase corresponds to the pattern retrieval region in the generalized Hopfield models barra2018phase under the critical capacity. In this region the student pattern tends directly to be aligned with one of the examples instead of learning their archetype, i.e. the teacher pattern. Unlike the sR regime, this is a phase where the machine is trying to learn by memorization and not by generalization alemanno2023hopfield .
5.3 The role of the S-RBM priors
@@
In Fig. (4) the phase diagrams of different S-RBMs with binary patterns ($\Omega_{\xi}=0$) are shown in the case of a binary T-RBM, i.e. $\Omega_{\hat{\xi}}=\Omega_{\hat{\tau}}=0$. One can observe the appearance of the eR phase at low temperature as soon as the student hidden unit starts to have a Gaussian tail, i.e. $\Omega_{\tau}>0$. The critical temperature at $\alpha=0$ is exactly $\Omega_{\tau}$. In the limit $\Omega_{\tau}\to 1$ this phase extends up to the critical capacity of the Hopfield model $\alpha=0.14$. As expected the critical size $\alpha_{c}$ of the triple point doesn’t change with the student priors while its temperature increases with $\Omega_{\tau}$. As a consequence the area of the sR phase expands in the direction of higher temperatures. This is typically an advantage for MC algorithms that samples from the posterior starting from high temperature like simulated annealing baldassi2018efficiency , hinton2012practical .
@@
In Fig. (5), we observe similar behavior for a dataset generated by a Hopfield like T-RBM, i.e. $\Omega_{\hat{\tau}}=1$, $\Omega_{\hat{\xi}}=0$. The different generation procedures significantly affect the number of data points required for the student to transition into the sR regime: as prescribed by Eq.(51), $\alpha_{c}$ is smaller than that of Fig.(4). One can in principle investigate the effect of changing $\Omega_{\xi}$, however we have verified that the phase diagram is not affected by that prior since it only enters in the self-overlap, and $d=1$ is always solution of Eqs. (10-13). Interestingly, the benefit of using Gaussian units—where at any given temperature, the sR region occurs at smaller $\alpha$—remains consistent regardless of the dataset properties. Specifically, it is not always advantageous for the student to align its hyperparameters with those of the teacher.
Monte Carlo simulations shown in Fig. (6) are in agreement with the replica analysis and illustrate how the minimal load $\alpha$ required to enter the sR phase changes with the choice of prior. Specifically, we fix the regularization strength for the student weights and set the temperature to $T=1.5$, then evaluate the learning magnetization for different values of $\Omega_{\tau}=0,0.5,1.0$. As expected, Gaussian hidden units facilitate signal retrieval, as also evident in the phase diagram of Fig. (4-5).
Figure 6: Agreement between theoretical replica analysis and Monte Carlo simulations. The magnetization computed from Eqs. (1013) is shown in blue, while the results from Monte Carlo simulations are shown in orange. The shaded area represents the range of values observed across runs. To compute the magnetization, we averaged over $N_{s}$ student pattern configurations $\xi$, sampled from the posterior distribution in Eq. (4). The figure compares the load $\alpha$ required to enter the signal retrieval (sR) phase at fixed student temperature $T=1.5$, for different choices of hidden unit priors.
5.4 The role of the T-RBM priors
@@
Figure 7: Effect of the teacher architecture ($\hat{\beta}=0.8$) on the student’s phase diagram. The signal’s prior is defined by $\Omega_{\hat{\xi}}=0$. The analyzed student is working with a Hopfield architecture: $\Omega_{\xi}=0$$\Omega_{\tau}=1$. The more $\hat{\tau}$ is gaussian the more easily the student retrieves the pattern $\bm{\hat{\xi}}$, in particular $P_{\text{triple}}$ decreases in both its coordinates.
Fig. (7) shows the effect of changing $\Omega_{\hat{\tau}}$. It is possible to appreciate the fact that the information present in the data increases notably the more the $\hat{\tau}$ prior becomes Gaussian. This can be seen by the reduced amount of generated samples we need to enter the sR regime (a similar effect of increasing $\hat{\beta}$ in alemanno2023hopfield ).
On the other hand, the critical temperature of the student at the triple point, $T_{c}$, is lowered. Nevertheless, if we fix an inference temperature, the student enters the sR regime more easily as $\Omega_{\hat{\tau}}\to 1$.
@@
Finally in Fig. (9) we explore the role of the dataset unit prior $\Omega_{s}$. The different values of $\Omega_{s}$ do not change the triple point $P_{c}$. As one can observe the only effect is the shift of the eR-SG transition line. This is related to the memorization capability of the S-RBM and is in agreement with the phase diagrams in barra2018phase .
5.5 Spherical regularization
In the previous sections we have seen that the term $z(\bm{\xi})^{-M}$ in the posterior plays the role of a regularization enforcing the Nishimori identity $d=1$. Without that regularization the model would be ill defined in the case of an architecture with both $\Omega_{\xi}$ and $\Omega_{\tau}$ different from zero. This is a typical problem with fully Gaussian disordered systems at low temperature barra2018phase , barra2012glassy , barra2014solvable , genovese2015legendre .
From a practical point of view is more convenient to replace that term directly with a spherical constraint $\delta\left(N-\sum_{i=1}^{N}\xi_{i}^{2}\right)$. The computation leading to the set of saddle point Eqs. (10-13) can be exactly followed with the only addition of a lagrange multiplier $\omega$ in the gaussian weight for the $\xi$ variable, i.e. its effective distribution reads as
$\hskip-28.45274ptP_{\Omega_{\xi}}(\xi)\exp\left(-\frac{\alpha\beta}{2}\left(\alpha\beta q_{\tau}-d_{\tau}+\omega\right)\xi^{2}+\xi\left(\alpha\sqrt{\hat{\beta}\beta}m_{\tau}\hat{\xi}+\sqrt{\alpha\beta q_{\tau}}z+\beta\Omega_{\tau}ps\right)\right)\,$, (52)
where $\omega$ has to be fixed to obtain $d=1$. Making a comparison with Eq.(14) one can see that the correct choice is $\omega=\langle\tau^{2}\rangle_{\tau}$. From the computational point of view it is more convenient to adjust the Lagrange multiplier during the solution of the saddle point equations. In Fig.(10) it is shown the phase diagram where the priors of both the weights and the student hidden variables have Gaussian tails. The weights have been regularized with a spherical constraint. One can compare these results with the right panel of Fig.(4), where $\Omega_{\xi}=0$. It is possible to see the presence of the eR phase, due to $\Omega_{\tau}=1$, that tends to quickly disappear as $\Omega_{\xi}\to 1$, in agreement with barra2018phase . It is also observed a bending of the sR region at low temperature that however is not particularly significant since that is the region where RSB corrections are expected. Finally, as expected, the triple point position doesn’t change.
Figure 10: Spherical regularization effect on the retrieval capability of the model (case with $\Omega_{\tau}=1$). Using binary data from a binary T-RBM ($\hat{\beta}=0.8$), it is possible to see that $\alpha_{c}$ and $T_{c}$ do not move. The only effect from the changes of $\Omega_{\xi}$ is the compression of the eReR phase.
6 Conclusions